A solution made by dissolving NaCl in water has a mole fraction of NaCl of 0.15....

  1. A solution made by dissolving NaCl in water has a mole fraction of NaCl of 0.15. What is the weight percent of NaCl in this solution? What is the molality of NaCl in this solution
Answer #1

thanks and hope you like the answer
Answer #2

To determine the weight percent of NaCl in the solution, we need to calculate the mass of NaCl in the solution and divide it by the total mass of the solution, then multiply by 100.

Given that the mole fraction of NaCl is 0.15, it means that for every 0.15 moles of NaCl, there are 1 mole of NaCl + 0.15 moles of water (since mole fraction is defined as moles of component / total moles of all components).

To calculate the weight percent of NaCl, we first need to determine the mass of NaCl. This can be done by multiplying the mole fraction of NaCl by the molar mass of NaCl.

The molar mass of NaCl is 58.44 g/mol (22.99 g/mol for Na + 35.45 g/mol for Cl).

Mass of NaCl = mole fraction of NaCl * molar mass of NaCl = 0.15 * 58.44 g/mol

To calculate the total mass of the solution, we need to consider the mass of NaCl and the mass of water.

The weight percent of NaCl can be calculated as: Weight percent of NaCl = (Mass of NaCl / Total mass of solution) * 100

Now let's calculate the weight percent of NaCl in the solution:

Mass of NaCl = 0.15 * 58.44 g/mol = 8.766 g

To calculate the total mass of the solution, we need to consider the mass of NaCl and the mass of water. The total mass of the solution is not given in the question, so we cannot provide a specific weight percent without this information.

As for the molality of NaCl in the solution, molality is defined as the moles of solute (NaCl) divided by the mass of the solvent (water) in kilograms.

Molality of NaCl = Moles of NaCl / Mass of water in kg

Since the mole fraction of NaCl is 0.15, it means that there are 0.15 moles of NaCl per mole of NaCl + 0.15 moles of water. Therefore, the moles of NaCl in the solution is also 0.15.

To calculate the mass of water in kg, we need to convert the mass of water from grams to kilograms.

Now let's calculate the molality of NaCl in the solution:

Mass of water = Total mass of solution - Mass of NaCl

Molality of NaCl = 0.15 moles / (Mass of water in kg)

Please provide the total mass of the solution so we can calculate the weight percent and molality more accurately.


answered by: mervetokaz